# Lyrics: Boosting Fine-grained Language-Vision Alignment via Semantic-aware Visual Objects

### 08 Dec 2023 v1

@@ 5.4 Ablation Study
To investigate the impact of the semantic-aware visual objects and fine-grained representation alignment on the performance of Lyrics, we conduct ablation studies during the pre-training stage. In w/o grounding quries, we solely utilize visual quries to align both spatial representation and image caption simultaneously, and then transfer the output embedding of visual quries to a LLM for instruction fine-tuning. In w/o spatial representation, the visual quries are required to learn local visual features from the image caption. Both scenarios necessitate corresponding adjustments to the respective pre-training tasks. As illustrated in Table 4, the information imperfection in vision-language representations alignment downgrades performance significantly across all datasets. The performance drop is more severe in datasets that involve spatial visual reasoning and generating (e.g. SciQA-Img, TextVQA and RefCOCOg), where the semantic-aware visual objects to the MQ-Former can guide visual features to attend to informative image regions. Furthermore, in w/o pre-training stage, MQ-Former relies solely on the vision-to-language generative learning to bridge the modality gap. We can find that without vision-language alignment stage, the grounding and visual quries are unable to extract informative feature according to the given image and text, thus almost losing the zero-shot understanding and generating capabilities.
@@ 5.5 Qualitative Results
We further provide the qualitative results for a complementary understanding of the instructed zero-shot image-to-text generation capability of Lyrics. As illustrated in Figure 4, we present the responses of Lyrics and various mainstream LVLMs, such as BLIP2 [25], InstructBLIP [11] and Shikra [6], under the same instruction and image inputs. In the absence of fine-grained visual signals, previous methods fail to accurately capture detailed information of visual objects involved in the instructions, necessary for performing counting, discerning colors, recognizing actions, and judging position. We observe that with the local visual features provided by the visual refiner, Lyrics can effectively avoid visual hallucinations and factual errors. It is reasonable to infer that Lyrics is capable of understanding and perceiving the visual objects within the image via two-stage fine-grained vision-language representation alignment and generative learning. For example, Lyrics can leverage spatial information and local visual feature provided by the visual refiner to perceive the number, color and motion of visual objects contained in images, as well as to explore the relative positions between the perceived visual objects. Therefore, in the first case, Lyrics can identify that there are four skiers in the image who are in a resting state, and indicate a flushed face of the skier and he wear dark red padded jacket and black trouser within specific spatial coordinates via referential dialogue. Moreover, Lyrics effectively inherit the commonsense understanding and logical reasoning capabilities of LLMs, which enable the model to deduce the symbolic meaning of text and the result of code execution. We also discover that our method impressively identifies objective entities, such as notable figures and locations, indicating that the knowledge of LLMs is effectively fed back to MQ-Former during the process of instruction fine-tuning.
5.6 Few-shot Learning on Vision-Language Tasks
@@
To further verify the efficient learning and knowledge generalization of Lyrics, we conduct in-context few-shot learning on the Flickr30k [40], OKVQA [31] and SciQA-Img [29] and TextVQA [48] datasets. Note that we adopt naive random sample to construct the few-shot exemplars, and report the averaged scores for five different seeds. As illustrated in Figure 5, with a similar number of parameters, Lyrics exhibits stable performance peak and upward trend across various visual-language tasks, even when compared to LVLMs with powerful backbones. The performance significant improvement under one in-context learning sample indicates that our model can better refer the instruction following knowledge within a short context window.
@@ 6 Conclusion
In this paper, we propose Lyrics, a two-stage fine-grained pre-training and instruction fine-tuning framework towards generalist large vision-language model. We introduce a visual refiner designed to extract abstract local visual feature and concrete spatial information, which is comprised of an image tagging module, an object detection module and a semantic segmentation module. We first connect the Multi-scale Querying Transformer (MQ-Former) to frozen image encoder and visual refiner and bootstrap vision-language representation alignment via multi-task pre-training. Then, we connect the MQ-Former to the LLMs to bootstrap vision-to-language generative learning via semantic-aware object. Lyrics achieves impressive results across various vision-language tasks, and demonstrates a real-world dialogue capability in commonsense-grounded image description, visual scene understanding and reasoning, referential dialogue.
Appendix A Dataset Details
In the first stage of pre-training, We primarily utilize a large-scale, weakly-labeled, web-crawled image-text pair dataset, which is composed of several publicly accessible data sources. As summarized in Table 5, We filter out image-text pairs with a similarity below 25 through an internally trained CLIP [42] model and use the cleaned data for pre-training. We clean these noisy data by several steps:
During instruction fine-tuning stage, we introduce high-quality and fine-grained vision-language annotation data. As summarized in Table 6, we divide the collected datasets into 7 tasks and only take the train dataset to fine-tune Lyrics simultaneously. We transform 24 datasets into the instruction tuning format and group them into 6 task categories. For text generation, we use the collected corpus Orca [35] and Alpaca [49] to recover the language generation ability. Image captioning data is COCO [7], cleaned CC3M and SBU. We use a mixture of publicly available data for the VQA task which includes GQA [20], VQAv2 [17], OCR-VQA [34], OK-VQA [31], AOK-VQA [46], DocVQA [33], TextVQA [48] and ChartQA [32]. For the referring expression comprehension and grounded captioning duality tasks, we uniformly construct training samples from GRIT [36], Visual Genome [23], RefCOCO [21], RefCOCO+ [30] and RefCOCOg [30]. In order to improve the instruction following and dialogue capabilities, we further collect various instruction fine-tuning datasets for real-world scenarios, including LLaVA [27], SVIT [62], M3IT [26], MULTIINSTRUCT [57] and LLaVAR [61].
Appendix B Summary of the Evaluation Benchmarks and Overfitting Authentication
We provide a detailed summary of the used evaluation benchmarks and corresponding metrics in Table 7. Since we aim to stimulate the model’s comprehension and generation capabilities in primary vision-language tasks, we believe that incorporating the training set of a specific dataset into the extensive training process will not result in the model overfitting to that dataset’s performance patterns. Consequently, we define evaluations that incorporate the training set into the instruction fine-tuning stage as few-shot evaluations, and those that do not include the training set in the instruction fine-tuning phase as zero-shot evaluations.
Therefore, we conduct an ablation study on the Nocaps [1], Flickr30k [40], VQAv2 [17], OKVQA [31], GQA [20], TextVQA [48] and RefCOCOg [30] datasets, in which we remove the train split of the corresponding dataset from the instruction fine-tuning stage, aiming to achieve zero-shot evaluation. As illustrated in Table 8, removing the train split of a specific dataset during the instruction fine-tuning stage does not significantly impact the model’s performance on that dataset’s test split. For example, by excluding the train split of the Visual Question Answering dataset VQAv2 from our vision-language generative learning, the model train in this manner exhibits only a 1.3% lower performance on the zero-shot evaluation of VQAv2 compared to the few-shot evaluation. Interestingly, this approach even yields slight performance improvements on some datasets. This phenomenon demonstrates that with adequate training on diverse datasets, LVLMs can develop robustness and generalization, preventing them from being overly specialized and overfitting to certain specific tasks and formats.
Dataset | Original | Cleaned | Remaining |
LAION-en | 2B | 200M | 10 |
LAION-COCO | 600M | 300M | 50 |
CC12M | 12M | 6M | 50 |
CC3M | 3M | 3M | 100 |
SBU | 1M | 0.9M | 90 |
COCO Caption | 0.6M | 0.6M | 100 |
Appendix C Instruction Templates for Instruction Fine-tuning and Zero-shot Inference
@@
Appendix D Additional Experiment Details on Multi-modal Leaderboards
To have a systematic understanding of the performance of Lyrics, we aim to leverage a quantitative metric in measuring the model’s instruction-following capability under the real-world conversation scenarios. Therefore, 2e evaluate the performance of Lyrics on three leaderboards: LLaVA [27], MME [15], and POPE, and compare it with mainstream generalist models such as BLIP2 [25], InstructBLIP [11], Shikra [6], and Qwen-VL [3].
The LLaVA leaderboard utilizes GPT-4 [37] to assess the quality of the responses generated by the LVLMs. Specifically, we randomly select 30 images from the COCO validation split, and generate three types of question (conversation, detailed description, complex reasoning) using the proposed data generation pipeline. LVLMs predicts the answers based on the question and input image. GPT-4 makes a reference prediction based on the question, ground-truth bounding boxes and captions, marking an upper bound of the teacher model. After obtaining the response from both models, we feed the question, visual information (in the format of captions and bounding boxes), and the generated responses from both assistants, to the GPT-4. GPT-4 evaluates the helpfulness, relevance, accuracy, and level of details of the responses from the assistants, and give an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. GPT-4 is also asked to provide a comprehensive explanation the evaluation, for us to better understand the models.
The MME leaderboard covers the examination of perception and cognition abilities across 14 subtasks. The perception includes the recognition of coarse-grained and fine-grained objects. The former identifies the existence, count, position, and color of objects. The latter recognizes movie posters, celebrities, scenes, landmarks, and artworks. The cognition includes commonsense reasoning, numerical calculation, text translation, and code reasoning. Since this leaderboard limits the output of model to two types (“yes” or “no”), it measure the metrics of accuracy and accuracy+ to validate the performance. The former is calculated based on each question, while the latter is based on each image where both of the two questions need to be answered correctly. We calculate the score of a subtask based on the sum of accuracy and accuracy+.
The POPE leaderboard formulates the evaluation of object hallucination as a binary classification task that prompts LVLMs to output “Yes” or “No”, e.g., “Is there a chair in the image?”. Questions whose answers are “Yes” can be directly built using ground-truth objects, while questions with the answer “No” can be built by sampling from negative objects. It employs three sampling strategies to verify whether LVLMs are prone to hallucinate specific objects, including: Random Sampling, which randomly samples the objects that do not exist in the image; Popular Sampling, which selects the top-k most frequent objects in the whole image dastaset that do not exist in the current image; Adversarial Sampling, which first ranks all objects according to their co-occurring frequencies with the ground-truth objects and then selects the top-k frequent ones that do not exist in the image.
As illustrated in Table 9, in various categories and objectives of visual-language instruction tasks, Lyrics surpasses other LVLMs by a large margin, from which we can draw the following conclusions: (1) In the LLaVA leaderboard, Lyrics demonstrates scores approaching those of GPT-4. This indicates that the model, in addition to inheriting the logical reasoning and language expression capabilities of LLMs, has successfully integrate substantial visual signals into its understanding. (2) The significant performance improvement of Lyrics on MME and POPE leaderboards indicates its ability to largely mitigate the effects of visual illusions. It is likely attributed to the semantic-aware visual objects introduced by the Visual Refiner, which enables the model to have a clearer understanding of concepts related to quantity, color, position, and morphology in the images. (3) Lyrics demonstrates robust instruction-following capability by adapting to variations in instructions, enabling it to execute visual-language tasks in various scenarios, such as visual-centric description, visual scene understanding and reasoning, commonsense-grounded image perception, referential dialogue.
Task | Dataset |
Image Captioning | COCO 2014, COCO 2017, CC3M, SBU |
Grounded Captioning | GRIT, Visual Genome, RefCOCO, RefCOCO+, RefCOCOg |
Visual Question Answering | GQA, VQAv2, OCR-VQA, OK-VQA, AOK-VQA, DocVQA, TextVQA, ChartQA |
Referring Expression Comprehension | GRIT, Visual Genome, RefCOCO, RefCOCO+, RefCOCOg |
Multi-modal Instruction | LLaVA, SVIT, M3IT, MULTIINSTRUCT, LLaVAR |
Pure-text Auto-regression | Orca, Alpaca |
Task | Dataset | Description | Split | Metric |
Image Captioning | Nocaps | Captioning of natural images | val | CIDEr() |
Flickr30K | Captioning of natural images | karpathy-test | CIDEr() | |
General VQA | VQAv2 | VQA on natural images | test-dev | VQA Score() |
OKVQA | VQA on natural images requiring outside knowledge | val | VQA Score() | |
GQA | VQA on scene understanding and reasoning | test-balanced | EM() | |
ScienceQA (Image Set) | Multi-choice VQA on a diverse set of science topics | test | Accuracy() | |
VizWiz | VQA on photos taken by people who are blind | test-dev | VQA Score() | |
Text-oriented VQA | TextVQA | VQA on natural images containing text | val | VQA Score() |
OCRVQA | VQA on images of book covers | test | EM() | |
Refer Expression Comprehension | RefCOCO | Refer grounding on natural images | val & testA & testB | Accuracy() |
RefCOCO+ | Refer grounding on natural images | val & testA & testB | Accuracy() | |
RefCOCOg | Refer grounding on natural images | val & test | Accuracy() |
model | Nocaps | Flickr30k | VQAv2 | OKVQA | GQA | TextVQA | RefCOCOg | |
val | test | |||||||
Lyrics | 126.8 | 85.4 | 81.2 | 58.2 | 62.4 | 69.4 | 87.23 | 88.26 |
w/o specific train split | 125.2 | 86.1 | 82.5 | 58.9 | 61.2 | 69.8 | 86.89 | 88.20 |
Model | LLaVA | MME | POPE | |||||||||
Perception | Cognition | Random |
BLIP-2 | 82.8 | 81.9 | 85.9 | 1293.84 | 290.00 | 85.98 | 82.29 | 74.20 | ||||
InstructBLIP | 77.4 | 83.8 | 89.3 | 1212.82 | 291.79 | 88.57 | 82.77 | 72.10 | ||||
Shikra | 72.8 | 80.6 | 83.9 | 1229.6 | 311.62 | 86.90 | 83.97 | 83.10 | ||||
Qwen-VL | 87.1 | 86.2 | 89.1 | 1487.58 | 360.71 | 89.18 | 85.24 | 81.72 | ||||
Lyrics | 88.3 | 89.8 | 93.4 | 1537.37 | 431.06 | 91.25 | 87.33 | 83.55 |

### 12 Apr 2024 v2

@@ 5.4 Ablation Study
In this section, we conduct ablation studies on the model architecture and training strategies to investigate the impact of semantic-aware visual objects and fine-grained representation alignment on the performance of Lyrics.
Model Architecture.
As illustrated in Table 5, we investigate the performance degradation of Lyrics on four benchmarks following the removal of the visual encoder (w/o ViT), object detection module (w/o ODM), semantic segmentation module (w/o SSM) and visual refiner (w/o VR). Generally, we use blank images to replace the original images as input for the specified module to represent the removal of the module. Firstly, the elimination of the ViT leads to performance declines across all tasks, attributable to the absence of global visual feature. Furthermore, We can observe that relying solely on either object detection or semantic segmentation modules results in insufficient local visual information. However, the concurrent removal of both leads to significant performance degradation across all datasets. Particularly, there is a 5.01% and 5.80% decrease in performance on the RefCOCOg and HallB benchmarks, respectively, demonstrating the significance of directly learning visual objects to grasp regional information.
Training Strategy.
In Figure (a), we present our investigation into the required quantity of high-quality image captions in the pre-training stage. Without vision-language alignment, Lyrics suffers from catastrophic forgetting where performance drastically degrades following instruction fine-tuning, and the model shows consistent gain with more pre-training data. Concurrently, with the increase in pre-training steps, the model can quickly adapt to straightforward VQA tasks (i.e., VQAv2) in instruction fine-tuning stage, whereas for complex visual reasoning and referential dialogue tasks (i.e., RefCOCOg, LLaVA and HallB), effective fitting requires sufficient visual-language alignment beforehand. In Figure (b), we conduct a comprehensive comparison of training the LLM, visual encoder and visual refiner with frozen and full parameters during the instruction fine-tuning stage, as well as partially training LLM with LoRA strategy. The training results demonstrate that freezing all parameters diminishes the fitting speed and performance, primarily due to the lack of self-adjustment limiting the LLM to receive visual signals. Furthermore, the fitting trends between using LoRA and full-parameter training are remarkably similar, confirming that adequate visual-language alignment ensures lightweight instruction fine-tuning is sufficient for the LVLM to master various vision-to-language dialogue scenarios.
@@ 5.5 Qualitative Results
We further provide the qualitative results for a complementary understanding of the instructed zero-shot image-to-text generation capability of Lyrics. As illustrated in Figure 5, we present the responses of Lyrics and various mainstream LVLMs, such as BLIP2 [29], InstructBLIP [11] and Shikra [6], under the same instruction and image inputs. In the absence of fine-grained visual signals that necessary for performing counting, discerning colors, recognizing actions and judging position, previous methods fail to accurately capture detailed information of visual objects involved in the instructions. We observe that with the visual refiner, Lyrics can effectively avoid visual hallucinations and factual errors. It is reasonable to infer that Lyrics can understand and perceive the visual objects within the image via two-stage fine-grained vision-language representation alignment and generative learning. For example, Lyrics can leverage spatial information and local visual feature provided by the visual refiner to perceive the number, color and motion of visual objects contained in images, as well as to explore the relative positions between the perceived visual objects. Therefore, in the first case, Lyrics can identify that there are four skiers in the image who are in a resting state, and indicate a flushed face of the skier and he wear dark red padded jacket and black trouser within specific spatial coordinates via referential dialogue. Moreover, Lyrics effectively inherit the commonsense understanding and logical reasoning capabilities of LLMs, which enable the model to deduce the symbolic meaning of text and the result of code execution. We also discover that our method impressively identifies objective entities, such as notable figures and locations, indicating that the knowledge of LLMs is effectively fed back to MQ-Former during the process of instruction fine-tuning. More examples are displayed in supplementary materials.
Dataset | Model | ||||
BLIP2 | Shikra | Qwen-VL | Lyrics | ||
VQAv2 | 0-shot | 65.0 | 77.4 | 78.2 | 81.2 |
1-shot | 67.4 | 79.7 | 80.5 | 85.2 | |
4-shot | 72.5 | 81.2 | 81.3 | 86.1 | |
RefCOCOg | 0-shot | - | 83.16 | 86.32 | 88.26 |
1-shot | - | 84.29 | 87.13 | 91.39 | |
4-shot | - | 85.02 | 87.88 | 91.36 | |
LLaVA | 0-shot | 38.1 | - | 67.7 | 76.9 |
1-shot | 40.6 | - | 68.2 | 79.1 | |
4-shot | 43.3 | - | 68.6 | 78.6 |
5.6 Few-shot Learning on VL Tasks
@@
To further verify the efficient learning and knowledge generalization of Lyrics, we conduct in-context few-shot learning on the VQAv2 [17], RefCOCOg (testset) [37] and LLaVA-Bench [31] datasets. Note that we adopt naive random sample to construct the few-shot exemplars, and report the averaged scores for five different seeds. As illustrated in Table 6, with a similar number of parameters, Lyrics exhibits stable performance peak and upward trend across various visual-language tasks, even when compared to LVLMs with powerful backbones. Notably, with one in-context learning sample, Lyrics achieves significant performance improvements across various tasks. Compared to existing methods, our model more effectively utilizes knowledge of instruction following within a short context window, demonstrating that diverse visual feature priors contribute to promoting autonomous segmentation of task frameworks by the model.
@@ 6 Conclusion
In this paper, we propose Lyrics, a two-stage fine-grained pre-training and instruction fine-tuning framework towards the generalist LVLM. We introduce a visual refiner designed to extract abstract local visual feature and concrete spatial information, which is comprised of an image tagging module, an object detection module and a semantic segmentation module. We first connect the Multi-scale Querying Transformer (MQ-Former) to frozen image encoder and visual refiner and bootstrap vision-language representation alignment via multi-task pre-training. Then, we connect the MQ-Former to the LLMs to bootstrap vision-to-language generative learning via semantic-aware object. Lyrics achieves impressive results across various vision-language tasks, and demonstrates a real-world dialogue capability in commonsense-grounded image description, visual scene understanding and reasoning, referential dialogue.
Appendix A Instruction Templates for Instruction Fine-tuning and Zero-shot Inference